Silver tarnishes in the presence of hydrogen sulfide and oxygen according to the below reaction.

How many liters of O2 at 100.˚C and 115 kPa would be needed to react with 10.00 grams of silver?

4Ag + 2H2S + O2 —> 2Ag2S + 2H2O


1
Expert's answer
2021-05-11T05:56:07-0400

By ideal gas equation -


PV = NRT.......1)


P =115 KPa = 115000 KPa


T=100"^{0}C" =373.15 k


R = 8.314 m"^{3} Pa \/k\/mol"


"N=\dfrac{m}{M}" m = given mass M= Molar mass


N "=\dfrac {10}{107}=" 0.093 moles


now putting all these value in equation ....1) we get-


115000"\times V =0.093\times 8.314\times 373.15"



V="\ 2.50\times 10^{-3}\ m^{3}"